3- Assume that the time required to pour a concrete floor for a structure (D) has a triangular distribution between 18 and 22 days with the most required time of 20 days.. Determine the following probabilities:

a. (5%)Prob (D > 21)

b. (5%) Prob (D ≤ 20)

c. (5%) Prob (D = 19)
Answer #1

here the triangular distribution has parameter a =18 , b=22 and m=20

a)

P(D<21)=1-(b-d)2/((b-a)*(b-m))=0.875

b)

P(D<20 ) =(d-a)2/((b-a)*(m-a)) =0.50

c)

P(D=19) =0 (since point probability over a continuous distribution is 0)